### v1

@@ 7 Conclusion and future outlook
Here, $({\cal D}_{1}\bar{\bm{e}})_{i}=(\bar{e}_{i}-\bar{e}_{i-1})/h$, $({\cal D}_{2}^{0}\bar{\bm{e}})_{i}=(\bar{e}_{i}-2\bar{e}_{i+1}+\bar{e}_{i+2})/h^{2}$, and $({\cal D}_{2}^{1}\bar{\bm{e}})_{i}=(\bar{e}_{i-1}-2\bar{e}_{i}+\bar{e}_{i+1})/h^{2}$. The vector $\bm{\nu}^{n+j}=\big{(}\nu_{1+1/2}^{n+j},\nu_{2+1/2}^{n+j},\ldots,\nu_{n_{x}+1/2}^{n+j}\big{)}\in\mathbb{R}^{n_{x}}$ is the vector of dissipation coefficients used in the LF numerical flux (12) at time $t_{n+j}+\delta t\vartheta$. Recall from Section SM4.1 that to develop a truncation error estimate for the linearized problem, its time dependence on $[t_{n},t_{n}+\delta t]$ is frozen at the point $t_{n}+\delta t\vartheta$, with $\vartheta\in[0,1]$. Also, we have defined the shorthands
$\gamma_{i}^{0,n}:=2\widetilde{b}_{i+1}^{0,n}+b_{i}^{0,n},\quad\gamma_{i}^{1,n}:=\widetilde{b}_{i+1}^{1,n}+2b_{i}^{1,n},$ (SM88)
@@
where, e.g., $b_{i}^{1,n}$ means the reconstruction coefficient $b_{i}^{1}$ from time $t_{n}+\delta t\vartheta$. Further, $\odot$” denotes element-wise product between vectors. Notice that the expressions in (SM87) are, by design, equivalent those in (SM69) when the general weights $\{\widetilde{b}^{\ell}\}$, $\{{b}^{\ell}\}$ are replaced with the optimal linear weights $\{\widetilde{d}^{\ell}\}$, $\{{d}^{\ell}\}$.
SM6.2 Approximating powers of the spatial discretization
@@
Recalling the form of the approximate truncation error operator ${\cal T}_{\rm ideal}^{n}$ in (SM76), we need also to estimate powers of the spatial discretization matrix. To approximate these we use a similar strategy to that used previously in Section SM5.2 in the context of standard linear spatial discretizations; see (SM74). Specifically, for the $k=1$ we now use an expression that is identical to that in (SM74):
@@
$\sum\limits_{j=0}^{m-1}\big{(}\widehat{L}^{n+j}\big{)}^{2}\approx{\cal D}_{1}\operatorname{diag}\bigg{(}\sum\limits_{j=0}^{m-1}\big{(}\bm{\alpha}^{n+j}\big{)}^{2}\bigg{)}{\cal D}_{1}^{\top}\quad\textrm{for}\quad k=1,$ (SM89)
@@
However, for $k=2$ we do something slightly different than (SM74). Namely, rather than using the FD discretization ${\cal D}_{3}$ that appears in (SM74), we use a FD discretization of the 3rd-derivative that makes use of the WENO weights, analogous to what we did in (SM86). This results in
$\begin{split}\sum\limits_{j=0}^{m-1}\big{(}\widehat{L}^{n+j}\big{)}^{4}&\approx-\frac{3}{4}\frac{1}{h}{\cal D}_{1}\bigg{[}\operatorname{diag}\bigg{(}\sum\limits_{j=0}^{m-1}\big{(}\bm{\alpha}^{n+j}\big{)}^{4}\odot\bm{\gamma}^{1,n+j}\bigg{)}{\cal D}_{2}^{1}\\ &\hskip 34.44434pt-\operatorname{diag}\bigg{(}\sum\limits_{j=0}^{m-1}\big{(}\bm{\alpha}^{n+j}\big{)}^{4}\odot\bm{\gamma}^{0,n+j}\bigg{)}{\cal D}_{2}^{0}\bigg{]}\quad\textrm{for}\quad k=2,\end{split}$ (SM90)
with elements of $\bm{\gamma}^{*,n+j}$ defined in (SM88).
SM6.3 Piecing it all together
Now we produce our approximation for ${\cal T}_{\rm ideal}^{n}$ based on (SM76). This requires the spatial discretization estimates from (SM87), and powers of the spatial discretization given in (SM89) and (SM90). Piecing these together results in the following:
${\cal T}_{\rm ideal}^{n}=\begin{cases}{\cal D}_{1}\operatorname{diag}\bigg{(}\sum\limits_{j=0}^{m-1}\bm{\beta}^{n+j}\bigg{)}{\cal D}_{1}^{\top},\quad k=1,\\[12.91663pt] {\cal D}_{1}\bigg{[}\operatorname{diag}\bigg{(}\sum\limits_{j=0}^{m-1}\bm{\beta}^{0,n+j}\bigg{)}{\cal D}_{2}^{0}-\operatorname{diag}\bigg{(}\sum\limits_{j=0}^{m-1}\bm{\beta}^{1,n+j}\bigg{)}{\cal D}_{2}^{1}\bigg{]},\quad k=2.\end{cases}$ (SM91)
@@
$\beta_{i}^{n}$ $:=\left[\widehat{e}_{\rm RK}\Big{(}\delta t\alpha^{n}_{i+1/2}\Big{)}^{2}+\frac{h\delta t}{2}\nu^{n}_{i+1/2}\right],$ (SM92)
$\beta_{i}^{*,n}$ $:=\frac{3}{4}\left[-\frac{\widehat{e}_{\rm RK}}{h}\Big{(}\delta t\alpha^{n}_{i+1/2}\Big{)}^{4}+\frac{h^{2}\delta t}{12}\nu^{n}_{i+1/2}\right]\gamma_{i}^{*,n}.$ (SM93)
Recall from Section SM4.1 that the ERK constants are $\widehat{e}_{\rm RK}=-1/2$ (for the forward Euler method (3), as used when $k=1$), and $\widehat{e}_{\rm RK}=-1/4!$ (for the 3rd-order ERK method (2), as used when $k=2$). We remark that, by design, when the general weights $\{\widetilde{b}^{\ell}\}$, $\{{b}^{\ell}\}$ are replaced with the optimal linear weights $\{\widetilde{d}^{\ell}\}$, $\{{d}^{\ell}\}$ the truncation error operators here reduce to those used in the case of a standard linear discretization given in (SM75) (modulo the fact that the wave-speeds and dissipation coefficients here are evaluated in terms of the linearization point).
SM7 Over-solving: Discretization and algebraic error for a Burgers problem
@@
$u_{0}(x)=\begin{cases}0,\quad&-1\leq x\leq-\tfrac{1}{2},\\ 1,\quad&-\tfrac{1}{2}<x<0,\\ 0,\quad&\hphantom{-}0\leq x.\end{cases}$ (SM94)
@@
$u_{\textrm{exact}}(x,t)=\begin{cases}u_{\textrm{pre-merge}}(x,t),\quad&0<t<1,\\ u_{\textrm{post-merge}}(x,t),\quad&1\leq t,\end{cases}$ (SM95)
@@
$u_{\textrm{pre-merge}}(x)=\begin{cases}0,\quad&-1\leq x\leq-\tfrac{1}{2},\\ \tfrac{x+1/2}{t},\quad&-\tfrac{1}{2}\leq x\leq x_{\textrm{front}}(t),\\ 1,\quad&x_{\textrm{front}}(t)\leq x\leq x_{\textrm{shock}}(t),\\ 0,\quad&x_{\textrm{shock}}<x,\end{cases}$ (SM96)
@@
$u_{\textrm{post-merge}}(x)=\begin{cases}0,\quad&-1\leq x\leq-\tfrac{1}{2},\\ \tfrac{x+1/2}{t},\quad&-\tfrac{1}{2}\leq x\leq x_{\textrm{shock}}(t),\\ 0,\quad&x_{\textrm{shock}}(t)<x,\end{cases}$ (SM97)
@@
$x_{\textrm{shock}}(t)=\begin{cases}\tfrac{t}{2},\quad&0<t\leq 1,\\ \sqrt{t}-\tfrac{1}{2},\quad&1\leq t,\end{cases},\quad x_{\textrm{front}}(t)=\begin{cases}t-\tfrac{1}{2},\quad&0<t\leq 1,\\ x_{\textrm{shock}}(t),\quad&1\leq t.\end{cases}$ (SM98)
@@
Figure SM7: Left: Discretized space-time solution of Burgers equation (B) obtained by time-stepping on a mesh with $n_{x}=1024$ FV cells, and using a LLF numerical flux, and 3rd-order WENO reconstructions. Right: Associated discretization error in $\log_{10}$ scale as computed via the exact solution (SM95). The discrete $L^{1}$- and $L^{\infty}$-norms of the discretization error are included in the plot’s title. To aid with visualization, cell-wise errors less than $10^{-8}$ have been masked to white.
We now consider a specific numerical example with $n_{x}=1024$ FV cells in space. A LLF numerical flux is used with 3rd-order WENO reconstructions. In the left panel of Fig. SM7 the space-time solution obtained by sequential time-stepping is shown. The right panel of Fig. SM7 shows the associated discretization error, as computed with the exact solution (SM95). More specifically, the discretization error at time $t_{n}$ in the $i$th FV cell is defined as $\bar{e}_{\textrm{disc},i}^{n}:=\bar{u}_{\textrm{exact},i}(t_{n})-\bar{u}_{i}^{n}$, where $\bar{u}_{i}^{n}$ is the time-stepping solution, and $\bar{u}_{\textrm{exact},i}(t_{n})$ is the cell average of the exact PDE solution (SM95). The discretization error is essentially zero in regions where the solution is constant. It is on the order of $0.2$ along the shock, and it is also relatively large along the front of the rarefaction wave before it merges into the shock. It is also somewhat large along the base of the rarefaction wave (i.e., around the line $x=-0.5$). Thus, unsurprisingly, the discretization error is greatest where the true solution lacks regularity.
Now we use our iterative parallel-in-time solver Algorithm 1 to solve the discretized problem. The solver setting are used as in the main paper, including a single MGRIT iteration to approximately solve the linearized problem and the use of nonlinear F-relaxation. A Newton linearization is used for the WENO weights. At the $k$th iteration iteration, define the algebraic error as $\bar{e}_{\textrm{alg},i}^{n}:=\bar{u}_{i}^{n}-\big{(}\bar{\bm{u}}_{k}^{n}\big{)}_{i}$, where $\bar{u}_{i}^{n}$ is again the exact solution of the discretized problem obtained via sequential time-stepping, and $\big{(}\bar{\bm{u}}_{k}^{n}\big{)}_{i}$ is the approximation at iteration $k$ of the discretized problem. The algebraic error for iterations $k=0,2,4,6$ is shown in Fig. SM8.131313Note that the error is shown for the $k$th iterate after nonlinear F-relaxation has been applied to it; that is, in the notation of Algorithm 1, the algebraic error is actually shown for $\widetilde{\bm{u}}_{k}$ rather than $\bar{\bm{u}}_{k}$.
@@
Figure SM8: Algebraic error in $\log_{10}$ scale of the iterates $\bar{\bm{u}}_{k}$ produced by Algorithm 1 for the discretized problem from Fig. SM7. The error is shown for iterations $k=0,2,4,6$, as indicated in the title of each plot. Note the discrete $L^{-1}$- and $L^{\infty}$-norms of each algebraic error is also shown in the title. To aid with visualization, cell-wise errors less than $10^{-8}$ have been masked to white. Note that the same colour scale is used for all plots.
@@
Considering the top left plot in Fig. SM8, the algebraic error for the initial iterate has a very similar structure to the discretization error shown in the right of Fig. SM7. Recall that the initial iterate is obtained by interpolating the exact discretized solution from an $n_{x}=512$ mesh. The key difference between the $n_{x}=1024$ solution compared to $n_{x}=512$ solution is that the distance that the shock and the base of the rarefaction wave are smeared across is halved. This is why the algebraic error is largest in these regions.
@@
As the iteration proceeds, the algebraic error decreases rapidly, both in $L^{1}$- and $L^{\infty}$-norm—the fact that the $L^{\infty}$-norm goes to zero shows that we really do converge point-wise to the sequential time-stepping solution. The algebraic error appears to be reduced much more slowly where the solution lacks regularity. We suspect this is because the linearization is least accurate in these regions where the gradient is large, or does not even exist. Ironically, the primary effect of later iterations (e.g., $k\geq 4$) is to reduce the algebraic error in non-smooth regions, despite the fact that this is where the discretization is least accurate (see right panel of Fig. SM7). [^14] This behavior of algebraic error rapidly reducing in smooth regions while more slowly decreasing in non-smooth regions is typical in our experience, even for problems with more complicated structure in smooth regions. For example, this occurs for the Buckley–Leverett test problem used in the paper (see the right of Fig. 2), and for several Burgers problems we have tested where a shock develops from smooth initial conditions. In other words, these later iterations do not result in an approximation that more accurately approximates the true PDE solution. In fact, the algebraic error at iteration $k=2$ is already a factor of 10 smaller than the discretization error both in $L^{1}$- and $L^{\infty}$-norm. Therefore, the $k=2$ iterate is very likely just as good an approximation to the true PDE solution as the exact time-stepping solution is. We note, however, at iteration $k=2$ that the algebraic error is not point-wise smaller than the discretization error everywhere in the domain. For example, there are neighbourhoods around the shock and base of the rarefaction wave where the discretization error is essentially machine zero, while the algebraic error is many orders of magnitude larger, even though the algebraic error is still relatively small in these regions, especially when compared to the discretization error in neighboring cells that are in or adjacent to the shock or the base of the rarefaction wave.

### v2

@@ Acknowledgments
Here, $({\cal D}_{1}\bar{\bm{e}})_{i}=(\bar{e}_{i}-\bar{e}_{i-1})/h$, $({\cal D}_{2}^{0}\bar{\bm{e}})_{i}=(\bar{e}_{i}-2\bar{e}_{i+1}+\bar{e}_{i+2})/h^{2}$, and $({\cal D}_{2}^{1}\bar{\bm{e}})_{i}=(\bar{e}_{i-1}-2\bar{e}_{i}+\bar{e}_{i+1})/h^{2}$. The vector $\bm{\nu}^{n+j}=\big(\nu_{1+1/2}^{n+j},\nu_{2+1/2}^{n+j},\ldots,\nu_{n_{x}+1/2}^{n+j}\big)\in\mathbb{R}^{n_{x}}$ is the vector of dissipation coefficients used in the LF numerical flux (10) at time $t_{n+j}+\delta t\vartheta$. Recall from Section SM6.1 that to develop a truncation error estimate for the linearized problem, its time dependence on $t_{n},t_{n}+\delta t]$ is frozen at the point $t_{n}+\delta t\vartheta$, with $\vartheta\in[0,1]$. Also, we have defined the shorthands
(SM98) $\gamma_{i}^{0,n}:=2\widetilde{b}_{i+1}^{0,n}+b_{i}^{0,n},\quad\gamma_{i}^{1,n}:=\widetilde{b}_{i+1}^{1,n}+2b_{i}^{1,n},$
@@
where, e.g., $b_{i}^{1,n}$ means the reconstruction coefficient $b_{i}^{1}$ from time $t_{n}+\delta t\vartheta$. Further, $\odot$” denotes element-wise product between vectors. Notice that the expressions in (SM97) are, by design, equivalent those in (SM79) when the general weights $\{\widetilde{b}^{\ell}\}$, $\{{b}^{\ell}\}$ are replaced with the optimal linear weights $\{\widetilde{d}^{\ell}\}$, $\{{d}^{\ell}\}$.
SM8.2 Approximating powers of the spatial discretization
@@
Recalling the form of the approximate truncation error operator ${\cal T}_{\rm ideal}^{n}$ in (SM86), we need also to estimate powers of the spatial discretization matrix. To approximate these we use a similar strategy to that used previously in Section SM7.2 in the context of standard linear spatial discretizations; see (SM84). Specifically, for the $k=1$ we now use an expression that is identical to that in (SM84):
@@
(SM99) $\sum\limits_{j=0}^{m-1}\big(\widehat{L}^{n+j}\big)^{2}\approx{\cal D}_{1}\operatorname{diag}\bigg(\sum\limits_{j=0}^{m-1}\big(\bm{\alpha}^{n+j}\big)^{2}\bigg){\cal D}_{1}^{\top}\quad\textrm{for}\quad k=1,$
@@
However, for $k=2$ we do something slightly different than (SM84). Namely, rather than using the FD discretization ${\cal D}_{3}$ that appears in (SM84), we use a FD discretization of the 3rd-derivative that makes use of the WENO weights, analogous to what we did in (SM96). This results in
(SM100) $\begin{split}\sum\limits_{j=0}^{m-1}\big(\widehat{L}^{n+j}\big)^{4}&\approx-\frac{3}{4}\frac{1}{h}{\cal D}_{1}\bigg[\operatorname{diag}\bigg(\sum\limits_{j=0}^{m-1}\big(\bm{\alpha}^{n+j}\big)^{4}\odot\bm{\gamma}^{1,n+j}\bigg){\cal D}_{2}^{1}\\ &\hskip 34.44434pt-\operatorname{diag}\bigg(\sum\limits_{j=0}^{m-1}\big(\bm{\alpha}^{n+j}\big)^{4}\odot\bm{\gamma}^{0,n+j}\bigg){\cal D}_{2}^{0}\bigg]\quad\textrm{for}\quad k=2,\end{split}$
with elements of $\bm{\gamma}^{*,n+j}$ defined in (SM98).
SM8.3 Piecing it all together
Now we produce our approximation for ${\cal T}_{\rm ideal}^{n}$ based on (SM86). This requires the spatial discretization estimates from (SM97), and powers of the spatial discretization given in (SM99) and (SM100). Piecing these together results in the following:
(SM101) ${\cal T}_{\rm ideal}^{n}=\begin{cases}{\cal D}_{1}\operatorname{diag}\bigg(\sum\limits_{j=0}^{m-1}\bm{\beta}^{n+j}\bigg){\cal D}_{1}^{\top},\quad k=1,\\[12.91663pt] {\cal D}_{1}\bigg[\operatorname{diag}\bigg(\sum\limits_{j=0}^{m-1}\bm{\beta}^{0,n+j}\bigg){\cal D}_{2}^{0}-\operatorname{diag}\bigg(\sum\limits_{j=0}^{m-1}\bm{\beta}^{1,n+j}\bigg){\cal D}_{2}^{1}\bigg],\quad k=2.\end{cases}$
@@
(SM102) $\beta_{i}^{n}$ $:=\left[\widehat{e}_{\rm RK}\Big(\delta t\alpha^{n}_{i+1/2}\Big)^{2}+\frac{h\delta t}{2}\nu^{n}_{i+1/2}\right],$
(SM103) $\beta_{i}^{*,n}$ $:=\frac{3}{4}\left[-\frac{\widehat{e}_{\rm RK}}{h}\Big(\delta t\alpha^{n}_{i+1/2}\Big)^{4}+\frac{h^{2}\delta t}{12}\nu^{n}_{i+1/2}\right]\gamma_{i}^{*,n}.$
Recall from Section SM6.1 that the ERK constants are $\widehat{e}_{\rm RK}=-1/2$ (for the forward Euler method (3), as used when $k=1$), and $\widehat{e}_{\rm RK}=-1/4!$ (for the 3rd-order ERK method (4), as used when $k=2$). We remark that, by design, when the general weights $\{\widetilde{b}^{\ell}\}$, $\{{b}^{\ell}\}$ are replaced with the optimal linear weights $\{\widetilde{d}^{\ell}\}$, $\{{d}^{\ell}\}$ the truncation error operators here reduce to those used in the case of a standard linear discretization given in (SM85) (modulo the fact that the wave-speeds and dissipation coefficients here are evaluated in terms of the linearization point).
SM9 Over-solving: Discretization and algebraic error for a Burgers problem
@@
(SM104) $u_{0}(x)=\begin{cases}0,\quad&-1\leq x\leq-\tfrac{1}{2},\\ 1,\quad&-\tfrac{1}{2}<x<0,\\ 0,\quad&\hphantom{-}0\leq x.\end{cases}$
@@
(SM105) $u_{\textrm{exact}}(x,t)=\begin{cases}u_{\textrm{pre-merge}}(x,t),\quad&0<t<1,\\ u_{\textrm{post-merge}}(x,t),\quad&1\leq t,\end{cases}$
@@
(SM106) $u_{\textrm{pre-merge}}(x)=\begin{cases}0,\quad&-1\leq x\leq-\tfrac{1}{2},\\ \tfrac{x+1/2}{t},\quad&-\tfrac{1}{2}\leq x\leq x_{\textrm{front}}(t),\\ 1,\quad&x_{\textrm{front}}(t)\leq x\leq x_{\textrm{shock}}(t),\\ 0,\quad&x_{\textrm{shock}}<x,\end{cases}$
@@
(SM107) $u_{\textrm{post-merge}}(x)=\begin{cases}0,\quad&-1\leq x\leq-\tfrac{1}{2},\\ \tfrac{x+1/2}{t},\quad&-\tfrac{1}{2}\leq x\leq x_{\textrm{shock}}(t),\\ 0,\quad&x_{\textrm{shock}}(t)<x,\end{cases}$
@@
(SM108) $x_{\textrm{shock}}(t)=\begin{cases}\tfrac{t}{2},\quad&0<t\leq 1,\\ \sqrt{t}-\tfrac{1}{2},\quad&1\leq t,\end{cases},\quad x_{\textrm{front}}(t)=\begin{cases}t-\tfrac{1}{2},\quad&0<t\leq 1,\\ x_{\textrm{shock}}(t),\quad&1\leq t.\end{cases}$
@@
Figure SM10: Left: Discretized space-time solution of Burgers equation (B) obtained by time-stepping on a mesh with $n_{x}=1024$ FV cells, and using a LLF numerical flux, and 3rd-order WENO reconstructions. Right: Associated discretization error in $\log_{10}$ scale as computed via the exact solution (SM105). The discrete $L^{1}$- and $L^{\infty}$-norms of the discretization error are included in the plot’s title. To aid with visualization, cell-wise errors less than $10^{-8}$ have been masked to white.
We now consider a specific numerical example with $n_{x}=1024$ FV cells in space. A LLF numerical flux is used with 3rd-order WENO reconstructions. In the left panel of Fig. SM10 the space-time solution obtained by sequential time-stepping is shown. The right panel of Fig. SM10 shows the associated discretization error, as computed with the exact solution (SM105). More specifically, the discretization error at time $t_{n}$ in the $i$th FV cell is defined as $\bar{e}_{\textrm{disc},i}^{n}:=\bar{u}_{\textrm{exact},i}(t_{n})-\bar{u}_{i}^{n}$, where $\bar{u}_{i}^{n}$ is the time-stepping solution, and $\bar{u}_{\textrm{exact},i}(t_{n})$ is the cell average of the exact PDE solution (SM105). The discretization error is essentially zero in regions where the solution is constant. It is on the order of $0.2$ along the shock, and it is also relatively large along the front of the rarefaction wave before it merges into the shock. It is also somewhat large along the base of the rarefaction wave (i.e., around the line x=0.5x=-0.5). Thus, unsurprisingly, the discretization error is greatest where the true solution lacks regularity.
Now we use our iterative parallel-in-time solver Algorithm 1 to solve the discretized problem. The solver setting are used as in the main paper, including a single MGRIT iteration to approximately solve the linearized problem and the use of nonlinear F-relaxation. A Newton linearization is used for the WENO weights. At the $k$th iteration iteration, define the algebraic error as $\bar{e}_{\textrm{alg},i}^{n}:=\bar{u}_{i}^{n}-\big(\bar{\bm{u}}_{k}^{n}\big)_{i}$, where $\bar{u}_{i}^{n}$ is again the exact solution of the discretized problem obtained via sequential time-stepping, and $\big(\bar{\bm{u}}_{k}^{n}\big)_{i}$ is the approximation at iteration $k$ of the discretized problem. The algebraic error for iterations $k=0,2,4,6$ is shown in Fig. SM11.131313Note that the error is shown for the $k$th iterate after nonlinear F-relaxation has been applied to it; that is, in the notation of Algorithm 1, the algebraic error is actually shown for $\widetilde{\bm{u}}_{k}$ rather than $\bar{\bm{u}}_{k}$.
@@
Figure SM11: Algebraic error in $\log_{10}$ scale of the iterates $\bar{\bm{u}}_{k}$ produced by Algorithm 1 for the discretized problem from Fig. SM10. The error is shown for iterations $k=0,2,4,6$, as indicated in the title of each plot. Note the discrete $L^{1}$- and $L^{\infty}$-norms of each algebraic error is also shown in the title. To aid with visualization, cell-wise errors less than $10^{-8}$ have been masked to white. Note that the same colour scale is used for all plots.
@@
Considering the top left plot in Fig. SM11, the algebraic error for the initial iterate has a very similar structure to the discretization error shown in the right of Fig. SM10. Recall that the initial iterate is obtained by interpolating the exact discretized solution from an $n_{x}=512$ mesh. The key difference between the $n_{x}=1024$ solution compared to $n_{x}=512$ solution is that the distance that the shock and the base of the rarefaction wave are smeared across is halved. This is why the algebraic error is largest in these regions.
@@
As the iteration proceeds, the algebraic error decreases rapidly, both in $L^{1}$- and $L^{\infty}$-norm—the fact that the $L^{\infty}$-norm goes to zero shows that we really do converge point-wise to the sequential time-stepping solution. The algebraic error appears to be reduced much more slowly where the solution lacks regularity. We suspect this is because the linearization is least accurate in these regions where the gradient is large, or does not even exist. Ironically, the primary effect of later iterations (e.g., $k\geq 4$) is to reduce the algebraic error in non-smooth regions, despite the fact that this is where the discretization is least accurate (see right panel of Fig. SM10). [^14] This behavior of algebraic error rapidly reducing in smooth regions while more slowly decreasing in non-smooth regions is typical in our experience, even for problems with more complicated structure in smooth regions. For example, this occurs for the Buckley–Leverett test problem used in the paper (see the right of Figure 2), and for several Burgers problems we have tested where a shock develops from smooth initial conditions. In other words, these later iterations do not result in an approximation that more accurately approximates the true PDE solution. In fact, the algebraic error at iteration $k=2$ is already a factor of 10 smaller than the discretization error both in $L^{1}$- and $L^{\infty}$-norm. Therefore, the $k=2$ iterate is very likely just as good an approximation to the true PDE solution as the exact time-stepping solution is. We note, however, at iteration $k=2$ that the algebraic error is not point-wise smaller than the discretization error everywhere in the domain. For example, there are neighbourhoods around the shock and base of the rarefaction wave where the discretization error is essentially machine zero, while the algebraic error is many orders of magnitude larger, even though the algebraic error is still relatively small in these regions, especially when compared to the discretization error in neighboring cells that are in or adjacent to the shock or the base of the rarefaction wave.